1. A rocket with a mass of 2.75 x^ 10 kg exerts a vertical force of 3.75 x ^10 N in the gases it expels. Determine your speed after 8.0s.
Answer #1

By Newton's second law:

F = ma

=> a = F/m

= (3.75*10^10 N)/(2.75*10^10 kg)

= 1.36 m/s^2

Given t = 8s

Use: v = at

= 1.36 * 8

= 10.9 m/s